### v1

# HoloTile RGB: Ultra-fast, Speckle-Free RGB Computer Generated Holography

Abstract
We demonstrate the first use of the HoloTile Computer Generated Holography (CGH) modality on multicolor targets. Taking advantage of the sub-hologram tiling and Point Spread Function (PSF) shaping of HoloTile allows for the reconstruction of high-fidelity, pseudo-digital RGB images, with well-defined output pixels, without the need for temporal averaging. For each wavelength, the target channels are scaled appropriately, using the same output pixel size. We employ a Stochastic Gradient Descent (SGD) hologram generation algorithm for each wavelength, and display them sequentially on a HoloEye GAEA 2.1 Spatial Light Modulator (SLM) in Color Field Sequential (CFS) phase modulation mode. As such, we get full 8-bit phase modulation at 60Hz for each wavelength. The reconstructions are projected onto a camera sensor where each RGB image is captured at once.
1 Introduction
Computer Generated Fourier holography (CGH) presents many advantages as a projection modality. While projection systems based on amplitude modulating digital micromirror displays (DMDs) are commonplace, they are accompanied by several downsides, including pattern-dependent light efficiency and single-plane focus. Using a phase-only spatial light modulator (SLM) in tandem with a Fourier transforming lens system allows for nearly 100% light efficiency in its reconstructions. In addition, the phase control of the SLM and the holographic encoding allows for dynamical shaping of light in both two or three dimensions. Several papers highlight the advantages of CGH in color reconstructions to generate beautiful reconstructions, however these typically suffer from the distinct speckle noise associated with CGH; a side-effect of both the coherent light involved and little to no control of the point spread function (PSF) of the optical system. Moreover, the calculation of high-resolution holograms necessary to take advantage of modern SLMs is typically too slow to be calculated in any real-time manner.
HoloTile [1, 2, 3, 4, 5, 6, 7, 8] is a CGH modality based on subhologram tiling and explicit PSF control. Tiling subholograms on the SLM has two distinct advantages; first, holograms are exponentially faster to calculate, and second, the tiling causes the reconstruction to only be defined in a very specific point grid, effectively separating the spatial frequency components. By adding a second, PSF controlling, hologram to the tiled hologram allows for complete control of the shape of each point in the reconstruction grid. Therefore, by shaping the PSF to be square, the reconstructions have effectively been discretized into a pseudo-digital output pixel grid. The well-defined output pixels do not overlap in the reconstruction plane and do not therefore contribute to the distinct speckle noise recognizable in conventional CGH. As opposed to other implementations of CGH [9], and particular RGB-CGH, it is not necessary to temporally average several holograms [10, 11, 12, 13, 14] in order to achieve beautiful reconstructions. Furthermore, since the sub-holograms are much smaller than the SLM resolution, calculation of HoloTile holograms can be performed at more than $60$Hz, allowing for real time calculation and display of scenes.
1
Input : Target amplitude $A_{t}$, iterations $T$, learning rate $lr=1e6$, parameter $\alpha=10$.
Output : Optimized hologram phase $\phi_{h}$.
2
[…]
4 $H=\text{FFT}(A_{t}\cdot e^{i\phi_{o}})$ // Compute initial complex hologram
5 $\phi_{h}=\angle H$ Extract hologram phase
6 for $i=1$ to $T$ do
7       $E_{r}\leftarrow\text{IFFT}(e^{i\phi_{h}})$ // Propagate back to object plane using inverse FFT
8       $A_{r}\leftarrow|E_{r}|$ // Compute reconstructed amplitude
9       $A_{r}\leftarrow A_{r}\cdot\sqrt{\frac{\sum A_{t}^{2}}{\sum A_{r}^{2}}}$ // Correct the energy
10       $\mathcal{L}\leftarrow\text{MSE}(A_{r}^{2},A_{t}^{2})$ // Calculate Mean Squared Error (MSE) loss
11
12      $\phi_{h}\leftarrow\phi_{h}-\text{lr}\cdot\nabla_{\phi_{h}}\mathcal{L}$ // Perform SGD update step on $\phi_{h}$
13
return $\phi_{h}$
Algorithm 1 SGD CGH Optimization
2 Theory
The process of generating HoloTile-based RGB computer generated holograms is relatively straightforward. Here, the process will be explained chronologically, i.e., the ordered computations that take place for each individual hologram generated for this paper.
2.1 Linear Color Space Conversion
Assuming an sRGB encoded target image is loaded into memory, a linear color space intensity version of the same image is calculated. This step ensures proper perception of the image when displayed using CGH, as no gamma correction is performed by the CGH reconstruction. The conversion from sRGB to linear color space is given by[15]
$I_{\text{lin}}=\begin{cases}\frac{I_{\text{sRGB}}}{12.92}&\text{for}\quad 0\leq I_{\text{sRGB}}\leq 0.04045\\ \left(\frac{I_{\text{sRGB}}+0.055}{1.055}\right)^{2.4}&\text{for}\quad 0.04045<I_{\text{sRGB}}\leq 1\end{cases}$ (1)
Figure 1: Lensed experimental targets captured directly in camera. Holograms are added a linear phase to filter out the zero order.
2.2 Output Grid Sampling and Alignment
The HoloTile modality defines an output grid in the reconstruction plane, with which the PSF shaping is convolved. The output “pixel” grid arises from the tiling of subholograms on the SLM. This grid can be expressed as
$A_{f}(x,y:t)\propto A_{\textrm{PSF}}(x,y,z;t)\otimes A_{\textrm{sub}}\left(x-\frac{m\lambda f}{\ell_{s}},y-\frac{n\lambda f}{\ell_{s}};t\right)$ (2)
where $\ell_{s}$ is the physical size of the subholograms, $\lambda$ is the illuminating wavelength, $f$ is the focal length of the Fourier transforming lens, $A_{\textrm{PSF}}$ is the Fourier transform of the PSF shaping hologram, and $A_{\textrm{sub}}$ is the Fourier transform of the subhologram. Due to the tiling operation, the reconstruction of the subhologram is sampled and thus only defined in specific locations defined by the spatial frequency comb array $\sum\limits_{m,n=-\infty}^{\infty}\delta\left(x-\frac{m\lambda f}{\ell_{s}},y-\frac{n\lambda f}{\ell_{s}}\right)$. To make sure that the three color channels of the reconstructions overlap exactly in the reconstruction plane, the output pixel interval must be identical:
$\frac{\lambda^{r}f}{\ell_{s}^{r}}=\frac{\lambda^{g}f}{\ell_{s}^{g}}=\frac{\lambda^{b}f}{\ell_{s}^{b}}=\ell_{p-out}$ (3)
As such, the size of the subholograms must be varied according to the illuminating wavelengths. Starting from defining $\ell_{p-out}$ to a desired output pixel interval, the subhologram sizes and, in turn, their resolutions $M^{\lambda}$, are calculated according to:
$\ell^{r}_{s}&=\frac{\lambda^{r}f}{\ell_{p-\text{out}}}&\quad\Rightarrow\quad&M^{r}=\left\lfloor\frac{\ell^{r}_{s}}{\ell_{p}}\right\rfloor\\ \ell^{g}_{s}&=\frac{\lambda^{g}}{\lambda^{r}}\ell^{r}_{s}&\quad\Rightarrow\quad&M^{g}=\left\lfloor\frac{\ell^{g}_{s}}{\ell_{p}}\right\rfloor\\ \ell^{b}_{s}&=\frac{\lambda^{b}}{\lambda^{r}}\ell^{r}_{s}&\quad\Rightarrow\quad&M^{b}=\left\lfloor\frac{\ell^{b}_{s}}{\ell_{p}}\right\rfloor\end{aligned}$ (4)
Figure 2: Experimental setup for capturing HoloTile RGB reconstructions shown in this publication. The spatial Fourier transforming convex lens is removed for the lensless configuration.
Where $\ell_{p}$ is the pixel pitch of the SLM. This will ensure the correct spacing of output pixels in the reconstruction plane. However, simply resizing the color channels to these resolutions will, necessarily, cause scaling issues, as the physical extent of the reconstructions are proportional to wavelength. To ensure that the reconstructions in the three color channels retain co-localization, all three color channels are first resized to the minimum subhologram resolution ($M^{b}$), and then zero-padded to the correct size according to its associated wavelength. The end result is then three targets, corresponding to the three color channels, RGB, of resolutions $M^{r}\times M^{r}$, $M^{g}\times M^{g}$, and $M^{b}\times M^{b}$, respectively.
2.3 CGH Optimization
To synthesize the sequential RGB hologram, each pre-processed color channel target is used to generate an associated phase-only subhologram. As HoloTile is agnostic towards the specific algorithm used for hologram generation, all reconstructions shown in this paper are calculated using Stochastic Gradient Descent (SGD) as presented in Algorithm 1. However, as shown in previous publications [2, 6], machine learning approaches or Gerchberg-Saxton-based (GS) algorithms are equally valid. The calculated wavelength-specific subhologram phases $\phi_{h}^{\lambda}$ are then tiled to the SLM resolution to create three tiled object holograms.
2.4 RGB PSF Shaping
To fill out the spatial frequency comb array that results from reconstructing only the tiled object holograms and create distinct output pixels, square PSF shaping is employed. As in the original HoloTile publication, square PSF shaping can be performed using an analytically computed phase profile, governed by $\beta$, a dimensionless parameter, given by
$\beta=\frac{2\sqrt{2\pi}R\ell_{p-out}}{f\lambda}$ (5)
where $R$ is the $1/e^{2}$ beam waist radius of the incident light on the SLM. Due to the wavelength dependence of the square beam shaping for any particular output pixel interval, a specific PSF shaping hologram is added to the tiled object hologram for each color channel to create the final HoloTile holograms. This ensures that the width of the output pixels are identical to the output pixel interval $\ell_{p-out}$.
3 Experimental Results
The laboratory setup used to generate the results of this paper is shown in Figure 2. A FISBA ReadyBeam Ind2 RGB Laser (wavelengths: 450nm, 520nm, and 638nm) is collimated and shone on a HoloEye GAEA 2.1 SLM. The reflected read-out beam is directed by a pellicle beam splitter through a convex lens performing a spatial Fourier transform, the result of which is captured directly on a CMOS chip (Canon EOS M6 Mark ii). Since the SLM is configured in CFS mode, each frame duration ($1/60s$) is divided into three subframes; one for each channel. The SLM is therefore cycling the color information at 180Hz.
Figure 3: Lensless experimental targets captured directly in camera. Holograms added a linear phase to filter out the zero order.
Four captured reconstructions directly from the camera are shown in Figure 1. To remove the zero-order diffraction spot and isolate the first order, a linear phase ramp was added to each color channel in the RGB holograms. In the reconstructions, the pseudo-digital output pixels are clearly seen. Furthermore, the output pixels are well-aligned in terms of size and color channel scaling. In addition, high-frequency speckle noise, as commonly seen in both monochrome and RGB CGH, is also absent due to the PSF shaping.
HoloTile can also function in a lensless configuration [6] by replacing the physical convex lens with an equivalent lens phase superimposed on the SLM. The lensless reconstructions as captured by the camera are shown in Figure 3. Here, the zeroth diffraction order is no longer a focused spot, but rather a large defocused white spot. Therefore, a linear phase ramp has been added to the lensless holograms as well in order to capture clean reconstructions. The lensless reconstructions also show the characteristic pseudo-digital output pixels, as well as good pixel alignment.
4 Conclusion
In this paper, we have demonstrated the first use of the HoloTile modality on multicolor targets. By leveraging the subhologram tiling and Point Spread Function shaping capabilities of HoloTile, we successfully reconstructed high-fidelity, pseudo-digital RGB images with well-defined output pixels, without the need for temporal averaging to reduce speckle. Each wavelength’s target channels were scaled appropriately, maintaining consistent output pixel sizes. Our approach utilized a Stochastic Gradient Descent hologram generation algorithm for each wavelength, displayed sequentially on a HoloEye GAEA 2.1 Spatial Light Modulator. This setup achieved full 8-bit phase modulation at 60Hz for each wavelength, with reconstructions captured directly on a camera sensor. The experimental results showcased the effectiveness of HoloTile in producing high-contrast reconstructions with superior homogeneity and regularity. The absence of high-frequency speckle noise, typically associated with conventional CGH, highlights the advantage of PSF shaping in our method. Additionally, the rapid hologram generation ($>$100x speed improvement) positions HoloTile as a promising technique for real-time applications in biophotonics, additive manufacturing, laser material processing, and advanced display technology.
This work has been supported by the Novo Nordisk Foundation, Denmark (Grand Challenge Program; NNF16OC0021948) and the Innovation Fund Denmark. We would like to express our gratitude to HoloEye Photonics AG for generously providing the HoloEye GAEA 2.1 Spatial Light Modulator used in this research.